# Globally-Constrained Decentralized Optimization with Variable Coupling

### v3

@@ III-C Decentralized Implementation
To obtain the update of $y_{i}^{k}$, let $\mathbf{d}^{k}=[d_{1}^{k},\ldots,d_{n}^{k}]$, where $d_{i}^{k}=\frac{\partial\mathcal{R}^{k}(\mathbf{y}^{k})}{\partial y_{i}}\!\in\!\mathbb{R}^{1\times(d_{i}+p)}$ $\forall i\in\mathcal{V}$ represents the Jacobian matrix of $\mathcal{R}^{k}(\mathbf{y})$ with respect to $y_{i}^{k}$. The expression of $d_{i}^{k}$ $\forall i\in\mathcal{V}$ directly follows from (12) and needs to compute $\frac{\partial\mathcal{R}^{k}(\mathbf{y}^{k})}{\partial y_{i}}$ by utilizing only elementary matrix operations, which is given
$\begin{split}d_{i}^{k}&=\Big{(}(\sum\limits_{j\in\mathcal{N}_{i}}\![P^{W}]_{ij}u_{j}^{k})^{T}-\frac{1}{\rho}(z_{i}^{k})^{T}\Big{)}B_{i}+[\tilde{d}_{i1}^{k},~{}\tilde{d}_{i2}^{k}],\end{split}$ (17)
@@
$\tilde{d}_{i1}^{k}$ $=\sum\limits_{j\in\mathcal{N}_{i}}\!\!\Big{(}\frac{\partial f_{j}(x^{k}_{\mathcal{N}_{j}})}{\partial x_{i}}+(q_{j}^{k}+g_{j}(x_{\mathcal{N}_{j}}^{k})-t_{j}^{k})^{T}\!\frac{\partial g_{j}(x^{k}_{\mathcal{N}_{j}})}{\partial x_{i}}\Big{)}$
$~{}~{}~{}~{}~{}+\frac{1}{\rho}\big{(}\!\sum\limits_{j\in\mathcal{N}_{i}}\!\!\!A_{ji}x_{i}^{k}\big{)}^{T}\sum\limits_{j\in\mathcal{N}_{i}}A_{ji}+\frac{1}{\rho}b_{i}^{T}\sum\limits_{j\in\mathcal{N}_{i}}A_{ji},$ (18)
$\begin{split}\tilde{d}_{i2}^{k}&=\frac{1}{\rho}(t_{i}^{k})^{T}-(q_{i}^{k}+g_{i}(x_{\mathcal{N}_{i}}^{k})-t_{i}^{k})^{T}.\end{split}$ (19)
The above equations (17)–(19) indicate that $d_{i}^{k}$ can be obtained by information interaction between neighboring nodes. Oracle 1 presents the decentralized calculation of $d_{i}^{k}$ $\forall i\in\mathcal{V}$.
@@
$y_{i}^{k+1}=\mathcal{P}_{Y_{i}}[y_{i}^{k}-\gamma\cdot(d_{i}^{k})^{T}]~{}~{}\forall i\!\in\!\mathcal{V}.$ (20)
@@
Algorithm 1 describes the detailed implementation of (20)–(23) taken by all the nodes. Given arbitrary $y_{i}^{0}\in Y_{i}$, $u_{i}^{0}\in\mathbb{R}^{p}$ $\forall i\in\mathcal{V}$, each node $i\!\in\!\mathcal{V}$ calculates $z_{i}^{0},q_{i}^{0}$ by collecting $u_{j}^{0},x_{j}^{0}$ from its every neighbor $j\in\mathcal{N}_{i}$. Then, at each $k\geq 0$, each node $i\in\!\!\mathcal{V}$ updates its primal variable $y_{i}^{k+1}$ according to (20), which requires node $i$ to calculate $d_{i}^{k}$ first. According to Oracle 1, each node $i$ needs to collect $x_{j}^{k}$, $u_{j}^{k}$, $A_{ji}$, and $\frac{\partial f_{j}(x^{k}_{\mathcal{N}_{j}})}{\partial x_{i}}+(q_{j}^{k}+g_{j}(x_{\mathcal{N}_{j}}^{k})-t_{j}^{k})^{T}\frac{\partial g_{j}(x^{k}_{\mathcal{N}_{j}})}{\partial x_{i}}\!$ from every neighbor node $j$ for obtaining $\tilde{d}_{i1}^{k},\tilde{d}_{i2}^{k}$ and then computes $d_{i}^{k}$ based on $\tilde{d}_{i1}^{k},\tilde{d}_{i2}^{k}$. The updates of $q_{i}^{k+1},u_{i}^{k+1}$ in (21), (22) can be realized by requesting $x_{j}^{k+1}$ from neighbor $j$. In addition, each node $i\in\mathcal{V}$ computes $z_{i}^{k+1}$ according to (23) by collecting $u_{j}^{k+1}$ from its neighbor $j$. Obviously, the updates (20)–(23) can be realized in a decentralized and inexpensive way.
TABLE 1: Oracle 1: Decentralized calculation of $d_{i}^{k}$ $\forall i\in\mathcal{V}$ $\forall k\geq 0$.
Upon receiving $x_{j}^{k}$, $u_{j}^{k}$ and $A_{ji}$ from every neighbor $j$, each node $i\in\mathcal{V}$
(a) calculates and sends $\frac{\partial f_{i}(x^{k}_{\mathcal{N}_{i}})}{\partial x_{j}}+(q_{i}^{k}+g_{i}(x_{\mathcal{N}_{i}}^{k})\!-t_{i}^{k})^{T}\frac{\partial g_{i}(x^{k}_{\mathcal{N}_{i}})}{\partial x_{j}}$ to its neighbor $j$;
(b) computes $\tilde{d}_{i1}^{k}$ according to (18);
(c) computes $\tilde{d}_{i2}^{k}$ according to (19);
(d) computes $d_{i}^{k}$ according to (17).
@@
2:  Each node $i\!\in\!\mathcal{V}$ sends $x_{i}^{0},\!\!\ u_{i}^{0}$ and $A_{ij}$ to every neighbor $j$ and sets $z_{i}^{0}\!=\rho\sum_{j\in\mathcal{N}_{i}}[P^{H}]_{ij}u_{j}^{0}$, $q_{i}^{0}=\max\{t_{i}^{0}-g_{i}(x_{\mathcal{N}_{i}}^{0}),\mathbf{0}_{p}\}$.
3:  for $k=0,1,2,\ldots$, each node $i\in\mathcal{V}$ do
4:     Compute $d_{i}^{k}$ according to Oracle 1;
5:     Update $y_{i}^{k+1}$ according to (20);
6:     Send $x_{i}^{k+1}$ to every neighbor $j$;
7:     Update $q_{i}^{k+1}$ according to (21);
8:     Update $u_{i}^{k+1}$ according to (22);
9:     Send $u_{i}^{k+1}$ to every neighbor $j$;
10:     Update $z_{i}^{k+1}$ according to (23).
11:  end for
III-C1 Special case of no variable coupling
@@
It is noteworthy that during initialization and each iteration $k\geq 0$, each node $i\in\mathcal{V}$ needs to communicate with its neighbors in lines 2, 4, 6, 9 of Algorithm 1. Nevertheless, if each $f_{i}$, $g_{i}$ only depends on $x_{i}$ and each $A_{i}$ only involves $x_{i}$, i.e., no variable coupling exists, the calculation of $d_{i}^{k}$ in Oracle 1 requires no information except $u_{j}^{k}$ from each neighbor $j$, which greatly simplifies the calculation of $d_{i}^{k}$. Moreover, in this case, updating $q_{i}^{k+1}$ in line 7 of Algorithm 1 does not rely on $x_{j}^{k}$ received from neighbor $j$, and updating $u_{i}^{k+1}$ in line 8 involves only $A_{i}$, no need of $A_{ji}$. The above simplified operators require each node $i\in\mathcal{V}$ to send only $u_{i}^{k}$ to its every neighbor $j$, and the transmission of $x_{i}^{0},A_{ij}$ in line 2, step (a) in Oracle 1, line 6 of Algorithm 1 can be canceled. Consequently, the communication and computational costs of Algorithm 1 can be significantly reduced. Additionally, in this case, all the nodes execute these simplified updates without disclosing their primal decisions, local objectives, and local constraint functions, thereby highly preserving information privacy.
@@ IV-A Convergence Results
TABLE 1: Table 1: Comparison with related methods in solving problems with globally-coupled inequality constraints yet with no variable coupling in local functions.
Algorithm compact domain smooth problem existence of equality constraint local optimization per iteration constant step-size convergence result
consensus-based primal-dual perturbation method [15] $\surd$ $\surd$ asymptotic
dual decomposition based method [16] $\surd$ $\surd$ asymptotic
RSDD [17] $\surd$ $\surd$ asymptotic
primal-dual gradient method [18] $\surd$ $\surd$ $\surd$ $\surd$ asymptotic
ALT [19] $\surd$ $\surd$ $\surd$ asymptotic
dual subgradient method [20] $\surd$ $\surd$ $\surd$ $O(\ln k/\sqrt{k})$
DPD-TV[21] $\surd$ $\surd$ $O(1/\ln k)$
$\text{DSA}_{2}$ [22] $\surd$ $\surd$ $O(1/\sqrt{k})$
C-SP-SG [23] $\surd$ $O(1/\sqrt{k})$
B-DPP [24] $\surd$ $\surd$ $O(1/\sqrt{k})$
IPLUX [25] $\surd$ $\surd$ $\surd$ $O(1/k)$
Algorithm 1 $\surd$ $\surd$ $\surd$ $\surd$ $O(1/k)$
@@
The existence of $\gamma,\rho>0$ under the parameter conditions in Theorem 1 is guaranteed and will be justified in Section IV-C. Additionally, although those theoretical parameter conditions involve the unavailable primal and dual optima, Section IV-C demonstrates that they can be satisfied based on accessible information only.
@@ IV-B Comparative Discussion
To the best of our knowledge, no prior methods can be directly applied to solve problem (1) in a decentralized way. Nevertheless, the algorithms in [15, 16, 17, 18, 20, 22, 21, 19, 23, 24, 25] are able to solve problems that admit globally-coupled inequality constraints yet no variable coupling, where each node’s local objective and constraint functions only depend on its own decision variable and are independent of its neighbors’ decisions. Here, we compare Algorithm 1 with those algorithms for solving such specialized forms of (1).
Observe from Table 1 that the algorithms in [15, 16, 17, 18, 19] do not have guaranteed convergence rates. The convergence rate results in [20, 22, 23, 21, 24] are weaker than the $O(1/k)$ rate for Algorithm 1, while they allow certain nonsmoothness in their problems. Although IPLUX [25] achieves $O(1/k)$ convergence rates in no need of a compact domain, it needs every node to solve a local convex optimization problem at each iteration (same for the algorithms in [24, 20, 19, 21, 17, 16, 22]), so that it has much higher computational complexity than Algorithm 1. Notably, only Algorithm 1 and [18, 20, 25, 19] consider problems with both inequality and equality constraints, while the others do not include equality constraints in their problem formulations. Although we may transform each equality constraint into two inequalities with reverse signs, this does not work for [15, 23, 21] because their methods require each inequality constraint to be strictly negative at a Slater point in order to bound the dual optimal set, and would significantly increase the variable dimensions and computational/communication complexities of the algorithms in [16, 17, 22, 24]. Furthermore, Algorithm 1 and [19, 18, 25] adopt constant step-sizes, whereas the remaining algorithms in Table 1 use diminishing step-sizes, which may lead to slow convergence in practice.

### v4

@@ III-C Decentralized Implementation
To obtain the update of $y_{i}^{k}$, let $\mathbf{d}^{k}=[(d_{1}^{k})^{T},\ldots,(d_{n}^{k})^{T}]^{T}$, where $d_{i}^{k}=\nabla_{y_{i}}\mathcal{R}^{k}(\mathbf{y}^{k})\!\in\!\mathbb{R}^{d_{i}+p}$ $\forall i\in\mathcal{V}$ represents the gradient of $\mathcal{R}^{k}(\mathbf{y})$ with respect to $y_{i}^{k}$. The expression of $d_{i}^{k}$ $\forall i\in\mathcal{V}$ directly follows from (12) and needs to compute $\nabla_{y_{i}}\mathcal{R}^{k}(\mathbf{y}^{k})$ by utilizing only elementary matrix operations, which is given
$d_{i}^{k}=B_{i}^{T}\Big(\!\sum\limits_{j\in\mathcal{N}_{i}}\![P^{W}]_{ij}u_{j}^{k}-\frac{1}{\rho}z_{i}^{k}\Big)+[(\tilde{d}_{i1}^{k})^{T},~(\tilde{d}_{i2}^{k})^{T}]^{T},$ (17)
@@
$\tilde{d}_{i1}^{k}$ $=\!\sum\limits_{j\in\mathcal{N}_{i}}\!\!\Big(\nabla_{x_{i}}f_{j}(x^{k}_{\mathcal{N}_{j}})+(\frac{\partial g_{j}(x^{k}_{\mathcal{N}_{j}})}{\partial x_{i}})^{T}(q_{j}^{k}+g_{j}(x_{\mathcal{N}_{j}}^{k})-t_{j}^{k})\!\Big)$
$~~~~~+\frac{1}{\rho}(\sum\limits_{j\in\mathcal{N}_{i}}A_{ji})^{T}\!\sum\limits_{j\in\mathcal{N}_{i}}\!\!\!A_{ji}x_{i}^{k}-\frac{1}{\rho}(\sum\limits_{j\in\mathcal{N}_{i}}A_{ji})^{T}b_{i},$ (18)
$\begin{split}\tilde{d}_{i2}^{k}&=\frac{1}{\rho}t_{i}^{k}-(q_{i}^{k}+g_{i}(x_{\mathcal{N}_{i}}^{k})-t_{i}^{k}).\end{split}$ (19)
The above equations (17)–(19) indicate that $d_{i}^{k}$ can be obtained by information interaction between neighboring nodes.
@@
$y_{i}^{k+1}=\mathcal{P}_{Y_{i}}[y_{i}^{k}-\gamma\cdot d_{i}^{k}]~~\forall i\!\in\!\mathcal{V}.$ (20)
@@
Algorithm 1 describes the detailed implementation of (20)–(23) taken by all the nodes. Given arbitrary $y_{i}^{0}\in Y_{i}$, $u_{i}^{0}\in\mathbb{R}^{p}$ $\forall i\in\mathcal{V}$, each node $i\!\in\!\mathcal{V}$ calculates $z_{i}^{0},q_{i}^{0}$ by collecting $u_{j}^{0},x_{j}^{0}$ from its every neighbor $j\in\mathcal{N}_{i}$. Then, at each $k\geq 0$, each node $i\in\!\!\mathcal{V}$ updates its primal variable $y_{i}^{k+1}$ according to (20), which requires node $i$ to calculate $d_{i}^{k}$ first. According to (17)–(19), each node $i$ needs to collect $x_{j}^{k}$, $u_{j}^{k}$, $A_{ji}$, and $\nabla_{x_{i}}f_{j}(x^{k}_{\mathcal{N}_{j}})+(\frac{\partial g_{j}(x^{k}_{\mathcal{N}_{j}})}{\partial x_{i}})^{T}(q_{j}^{k}+g_{j}(x_{\mathcal{N}_{j}}^{k})-t_{j}^{k})$ from every neighbor node $j$ for obtaining $\tilde{d}_{i1}^{k},\tilde{d}_{i2}^{k}$ and then computes $d_{i}^{k}$ based on them. The updates of $q_{i}^{k+1},u_{i}^{k+1}$ in (21), (22) can be realized by requesting $x_{j}^{k+1}$ from neighbor $j$. In addition, each node $i\in\mathcal{V}$ computes $z_{i}^{k+1}$ according to (23) by collecting $u_{j}^{k+1}$ from its neighbor $j$. Obviously, the updates (20)–(23) can be realized in a decentralized and inexpensive way.
@@
2: Each node $i\!\in\!\mathcal{V}$ sends $x_{i}^{0}$ and $A_{ij}$ to every neighbor $j$ and sets $q_{i}^{0}=\max\{t_{i}^{0}-g_{i}(x_{\mathcal{N}_{i}}^{0}),\mathbf{0}_{p}\}$.
3: Each node $i\!\in\!\mathcal{V}$ computes $\nabla_{x_{j}}f_{i}(x^{0}_{\mathcal{N}_{i}})+(\frac{\partial g_{i}(x^{0}_{\mathcal{N}_{i}})}{\partial x_{j}})^{T}(q_{i}^{0}+g_{i}(x_{\mathcal{N}_{i}}^{0})\!-t_{i}^{0})$, sends it with $u_{i}^{0}$ to every neighbor $j$, and sets $z_{i}^{0}\!=\rho\sum_{j\in\mathcal{N}_{i}}[P^{H}]_{ij}u_{j}^{0}$.
4:for $k=0,1,2,\ldots$, each node $i\in\mathcal{V}$ do
5:  Compute $d_{i}^{k}$ according to (17)–(19);
6:  Update $y_{i}^{k+1}$ according to (20);
7:  Send $x_{i}^{k+1}$ to every neighbor $j$;
8:  Update $q_{i}^{k+1}$ according to (21);
9:  Update $u_{i}^{k+1}$ according to (22);
10:  Compute $\nabla_{x_{j}}f_{i}(x^{k+1}_{\mathcal{N}_{i}})+(\frac{\partial g_{i}(x^{k+1}_{\mathcal{N}_{i}})}{\partial x_{j}})^{T}(q_{i}^{k+1}+g_{i}(x_{\mathcal{N}_{i}}^{k+1})\!-t_{i}^{k+1})$ and sent it with $u_{i}^{k+1}$ to every neighbor $j$;
11:  Update $z_{i}^{k+1}$ according to (23).
12:end for
@@
Special case of no variable coupling: It is noteworthy that at each iteration, each node $i\in\mathcal{V}$ needs to communicate with its neighbors twice. Nevertheless, if each $f_{i}$, $g_{i}$ only depends on $x_{i}$ and each $A_{i}$ only involves $x_{i}$, i.e., no variable coupling exists, the calculation of $d_{i}^{k}$ in (17) requires no information except $u_{j}^{k}$ from each neighbor $j$, which greatly simplifies the calculation of $d_{i}^{k}$. Moreover, in this case, updating $q_{i}^{k+1}$ in Line 8 of Algorithm 1 does not rely on $x_{j}^{k}$ received from neighbor $j$, and updating $u_{i}^{k+1}$ in Line 9 involves only $A_{i}$ and does not need $A_{ji}$. The above simplified operations require each node $i\in\mathcal{V}$ to send only $u_{i}^{k}$ to its every neighbor $j$, and the transmission of $x_{i}^{0},A_{ij}$ in Line 2, $x_{i}^{k+1}$ in Line 7, and $\nabla_{x_{j}}f_{i}(x^{k}_{\mathcal{N}_{i}})+(\frac{\partial g_{i}(x^{k}_{\mathcal{N}_{i}})}{\partial x_{j}})^{T}(q_{i}^{k}+g_{i}(x_{\mathcal{N}_{i}}^{k})\!-t_{i}^{k})\forall k\geq 0$ in Line 3 and Line 10 of Algorithm 1 can be removed. Consequently, the communication and computational costs of Algorithm 1 can be significantly reduced. Additionally, in this case, all the nodes do not have to share their primal decisions, the gradient of local cost functions, the values of local constraint functions, and their Jacobian matrices, thereby reducing the exposure of critical local information.
Remark 3.
Here, we discuss the differences and connections between Algorithm 1 and the existing methods applicable to various variable coupling [todescato2020partition, pu2016quantization, HuJH2018, XiaoYY2017, wang2024distributed, Alghunaim2019, Mota2014]. By letting each node maintain local variables and the copies of its neighbors’ variables, [HuJH2018] transforms its locally-coupled problem into an equality-constrained problem and proposes several algorithms based on operator splitting techniques to solve this equivalent problem. A similar strategy is adopted in [XiaoYY2017] to address convex feasibility problems. The methods in [Mota2014, Alghunaim2019] are able to tackle local functions involving non-neighbors’ variables (yet they cannot address globally-coupled constraints). In the special case where the local functions only depend on neighboring variables, they propose different cluster partition approaches, leading to the extended ADMM in [Mota2014] and the coupled diffusion strategy in [Alghunaim2019], where each node also needs to update its local variable and the copies of its neighbors’ variables. Clearly, the methods in [HuJH2018, XiaoYY2017, Alghunaim2019, Mota2014], incur higher memory and communication costs compared to Algorithm 1, especially for densely connected networks.
Like Algorithm 1, the methods in [todescato2020partition, pu2016quantization, wang2024distributed] allow each node to maintain its own decision variables only. Notably, in the absence of constraints or with only local set constraints, Algorithm 1 is equivalent to the synchronous gradient-based method in [todescato2020partition] and the unquantized version of the algorithm in [pu2016quantization]. Furthermore, by setting $W=I$, $\mathbf{z}^{k}=\mathbf{0}$, and $\mathbf{y}^{k}=\mathbf{x}^{k}$, our algorithm reduces to the method in [wang2024distributed] which only considers variable coupling under locally-coupled constraints. In contrast with [todescato2020partition, pu2016quantization, HuJH2018, XiaoYY2017, wang2024distributed, Alghunaim2019, Mota2014], our algorithm is able to simultaneously handle globally-coupled nonlinear inequality and linear equality constraints, making it applicable to a broader class of problems.
@@ IV-A Convergence Results
The inequality (24) specifies the conditions on the step-size $\gamma$ and the parameter $\rho$ required for Algorithm 1 to achieve the convergence rates in Theorem 1 and Theorem 2. The existence of positive $\gamma$ and $\rho$ satisfying (24) will be justified in Section IV-C. Additionally, although those theoretical parameter conditions involve the unavailable primal and dual optima, Section IV-C demonstrates that they can be satisfied based on accessible information only.
@@
Remark 4.
The network connectivity influences the convergence of constraint violations and objective error via the smallest nonzero eigenvalue of $H$, i.e., $\lambda_{\min}(H)$. This is because $\|\mathbf{z}^{0}-\mathbf{z}^{\star}\|_{H^{\dagger}}^{2}$, which determines the bounds on the right-hand side of (28)–(30), is bounded by $\frac{1}{\lambda_{\min}(H)}\|\mathbf{v}\|^{2}$, where $\mathbf{v}\in\mathbb{R}^{n(m+p)}$ such that $\mathbf{z}^{0}-\mathbf{z}^{\star}=H\mathbf{v}$. Therefore, larger $\lambda_{\min}(H)$, often indicating denser network connectivity, leads to faster convergence. On the other hand, denser network connectivity increases the computation and communication costs during each iteration. For example, adding a neighbor $\ell$ to node $i$ requires computing additional $d_{\ell}$ variables and requires transmitting $d_{\ell}+d_{i}+m+p$ additional real numbers.
TABLE 1: Comparison with related methods in solving problems with globally-coupled inequality constraints yet with no variable coupling in local functions.
Algorithm compact domain smooth problem existence of equality constraint main computational cost constant step-size convergence result
consensus-based primal-dual perturbation method [THChang2014] $\surd$ $\surd$ projection asymptotic
dual decomposition based method [Falsone2017] $\surd$ subproblem asymptotic
RSDD [Notarnicola2019] $\surd$ subproblem asymptotic
primal-dual gradient method [LiangS2019] $\surd$ $\surd$ $\surd$ projection $\surd$ asymptotic
ALT [falsone2023augmented] $\surd$ subproblem $\surd$ asymptotic
dual subgradient method [LiangS2019a] $\surd$ $\surd$ subproblem $O(\ln k/\sqrt{k})$
DPD-TV[camisa2021distributed] $\surd$ subproblem $O(1/\ln k)$
$\text{DSA}_{2}$ [LiuC2020] $\surd$ subproblem $O(1/\sqrt{k})$
C-SP-SG [Mateos2016] $\surd$ projection $O(1/\sqrt{k})$
B-DPP [wang2023distributed] $\surd$ subproblem $O(1/\sqrt{k})$
IPLUX [wu2022distributed] $\surd$ subproblem $\surd$ $O(1/k)$
Algorithm 1 $\surd$ $\surd$ $\surd$ projection $\surd$ $O(1/k)$
@@ IV-B Comparative Discussion
To the best of our knowledge, no prior methods can be directly applied to solve problem (1) in a decentralized way. Nevertheless, the algorithms in [THChang2014, Falsone2017, Notarnicola2019, LiangS2019, LiangS2019a, LiuC2020, camisa2021distributed, falsone2023augmented, Mateos2016, wang2023distributed, wu2022distributed] are able to solve problems that admit globally-coupled inequality constraints yet no variable coupling, where each node’s local objective and constraint functions only depend on its own decision variable and are independent of its neighbors’ decisions. Here, we compare Algorithm 1 with those algorithms for solving such specialized forms of (1).
Observe from TABLE 1 that the algorithms in [THChang2014, Falsone2017, Notarnicola2019, LiangS2019, falsone2023augmented] do not have guaranteed convergence rates. The convergence rate results in [LiangS2019a, LiuC2020, Mateos2016, camisa2021distributed, wang2023distributed] are weaker than the $O(1/k)$ rate for Algorithm 1, while they allow certain nonsmoothness in their problems. Although IPLUX [wu2022distributed] achieves $O(1/k)$ convergence rates in no need of a compact domain, it needs every node to solve a local convex optimization subproblem at each iteration (same for the algorithms in [wang2023distributed, LiangS2019a, falsone2023augmented, camisa2021distributed, Notarnicola2019, Falsone2017, LiuC2020]), so that it has higher computational complexity than Algorithm 1, which adopts a projection operation, i.e., minimizing a positive definite quadratic function subject to the same convex constraints. Notably, only Algorithm 1 and [LiangS2019, LiangS2019a, wu2022distributed, falsone2023augmented] consider problems with both inequality and equality constraints, while the others do not include equality constraints in their problem formulations. Although we may transform each equality constraint into two inequalities with reverse signs, this does not work for [THChang2014, Mateos2016, camisa2021distributed] because their methods require each inequality constraint to be strictly negative at a Slater point in order to bound the dual optimal set, and would significantly increase the variable dimensions and computational/communication complexities of the algorithms in [Falsone2017, Notarnicola2019, LiuC2020, wang2023distributed]. Furthermore, Algorithm 1 and [falsone2023augmented, LiangS2019, wu2022distributed] adopt constant step-sizes, whereas the remaining algorithms in TABLE 1 use diminishing step-sizes, which may lead to slow convergence in practice.
Furthermore, we would like to point out that the methods in [15]–[25] cannot be extended to tackle problem (1) with variable coupling in a straightforward way. This is because the distributed methods in [LiangS2019a, falsone2023augmented, camisa2021distributed, Notarnicola2019, Falsone2017, LiuC2020] inherently rely on the separability of the Lagrangian dual problem, which, however, cannot be decomposed into $n$ local optimization problems with $x_{i}$ being the sole optimization variable when the variable coupling in $f_{i}$ and $g_{i}$ exists. The works [wang2023distributed, wu2022distributed] update local primal variables by solving convex optimization problems that involve $f_{i}$ and $g_{i}$ at each iteration, which also cannot be implemented in a decentralized manner due to the variable coupling structure. The projection-based methods [THChang2014, LiangS2019, Mateos2016] either fail to handle the equality constraints [THChang2014, Mateos2016] as we stated before, or cannot ensure convergence [LiangS2019] when addressing problem (1).